# Implicit Personalization in Language Models: A Systematic Study

### v1

@@ Abstract
Implicit Personalization (IP) is a phenomenon of language models inferring a user’s background from the implicit cues in the input prompts and tailoring the response based on this inference. While previous work has touched upon various instances of this problem, there lacks a unified framework to study this behavior. This work systematically studies IP through a rigorous mathematical formulation, a multi-perspective moral reasoning framework, and a set of case studies. Our theoretical foundation for IP relies on a structural causal model and introduces a novel method, indirect intervention, to estimate the causal effect of a mediator variable that cannot be directly intervened upon. Beyond the technical approach, we also introduce a set of moral reasoning principles based on three schools of moral philosophy to study when IP may or may not be ethically appropriate. Equipped with both mathematical and ethical insights, we present three diverse case studies illustrating the varied nature of the IP problem and offer recommendations for future research. [^1] Our code and data are at https://github.com/jiarui-liu/IP.
@@ 1 Introduction
Inspired by this example, we propose the concept of Implicit Personalization (IP). Grounded in a structural causal model (SCM; Peters et al., 2017; Pearl, 2009), we define IP as a process that first infers a user’s background from the way a question is posed, and then tailors the response to fit this background, as in Figure 1. While many studies have separately explored different aspects of this problem (Flek, 2020; Raharjana et al., 2021), we still lack a community-wide standardized framework to study these phenomena. The absence of a common framework leads to divergent perspectives: some studies view it positively, suggesting that incorporating inferred user demographics can enhance NLP performance by personalized responses (Hovy, 2015; Benton et al., 2016; Sasaki et al., 2018; Salemi et al., 2024; Chen et al., 2023), whereas others criticize it negatively for introducing biases in model responses towards underrepresented groups (Bolukbasi et al., 2016; Garg et al., 2018; Arora et al., 2023; Das et al., 2023; He et al., 2024), or for fostering flattery to satisfy users regardless of the accuracy of the information provided (Sharma et al., 2023; Wei et al., 2023; Wang et al., 2023).
@@
  1. Q1.
    What is IP? – for which we provide a rigorous mathematical formalization (Section 2).
  2. Q2.
    How to detect IP in large language models (LLMs)? – for which we provide a rigorous mathematical formulation and a set of case studies (Sections 2, 4, 5, 6 and 7).
  3. Q3.
    What are the moral implications of IP? – for which we propose a framework for ethical engagement (Section 3).
  4. Q4.
    How to improve future models to ensure their IP behavior aligns with ethical standards? – for which we provide a list of suggestions for developers and the community (Section 8).
By answering the above questions, our work contributes a “full-stack” systematic study on IP: In the mathematical framework, we ground IP in an SCM Peters et al. (2017); Pearl (2009), and then propose an indirect intervention method to test the causal effect in the LLM-specific, diamond-shaped causal graph in Figure 1 with un-intervenable mediators (see technical deductions in Section 2). After the technical formulation of IP, we we provide a moral reasoning framework (Section 3), which connects the ethical considerations of IP to major schools of moral philosophy, including consequentialism Mill (2016); Parfit (1987), deontology Kant and Schneewind (2002); Ross (2002), and contractualism Rawls (2017); Scanlon (2000).
To illustrate the usefulness of our theoretical formulations, we present three diverse case studies that feature different natures of the IP problem (Section 4): (1) cultural adaptation, where IP is a desired model behavior (Section 5), (2) education disparity, where IP is unethical (Section 6), and (3) echo chamber, which has mixed implications for IP (Section 7). Finally, we conclude with recommendations for future research and an outlook for the community (Section 8).
@@ 2 A Mathematical Framework for IP
2.1 Notations
In general, an NLP system has the functional behavior $f:\bm{x}\mapsto\bm{y}$, where the user input the text $\bm{x}$, and the model generates a response $\bm{y}$, as in Figure 1. Formally, IP is a sub-process within this functional behavior. In our running example, where $\bm{x}=$“What color is a football?”, the model response $\bm{y}$ should be “Brown” if IP takes place, whereas the general answer would mention both possibilities, e.g., “An American football is Brown, and a soccer ball is usually black and white.”
We model this process of IP with the structural causal model $\mathcal{M}$ (SCM; Peters et al., 2017; Pearl, 2009) in Figure 1. For simplicity, we assume that the information contained in the input $\bm{x}$ can be fully represented by a tuple $(\bm{s},b)$, as the behavior of interest in this paper is the model reaction to the (implicitly) inferred user background. In this SCM, the model parses the input $\bm{x}$ and then generates a response $\bm{y}$. When parsing the input $\bm{x}$, the model grasps its semantics $\bm{s}$, namely asking about the color of a ball in the sport called football. The semantics $\bm{s}$ is further used to generate response $\bm{y}$. In the meantime, the model can choose to infer the user background $b$ (in this case, an American English speaker) from a categorical set of backgrounds $\mathcal{B}$. At the end, the model can choose to generate an answer only in response to the semantics $\bm{s}$ (i.e., without IP), or customize its answer to target at the user background $b$ (i.e., with IP).
2.2 Problem Formulation
We focus on the question: “Does IP take place in LLMs?” In terms of the SCM $\mathcal{M}$ framework, we reformulate this question as “Does user background ${B}$ have a causal effect on LLM’s response $\bm{Y}$?” This causal effect Peters et al. (2017) is defined as
@@ 2.2 Problem Formulation
In the SCM $\mathcal{M}$, there is a causal effect from $B$ to $\mathbf{Y}$ if there exist $b_{0},b_{1}\in\mathcal{B}$, such that
$P_{\bm{Y}}^{\mathcal{M}:\text{do}(B:=b_{0})}\neq P_{\bm{Y}}^{\mathcal{M}:\text{do}(B:=b_{1})}.$ (1)
The intuition behind this is we first intervene on $B$ by setting it to different values $b_{0}$ and $b_{1}$, and then compare whether the intervened probability distributions of $\bm{Y}$ are identical. If not, then it implies that the LLM performs IP to generate different response $\bm{Y}$ for different background $B$. Using this definition, we perform the following deduction:
@@
 $&\exists b_{0},b_{1}\in\mathcal{B},\text{s.t. }P_{\bm{Y}}^{\mathcal{M}:\operatorname*{do}(B:=b_{0})}\\ &\quad\quad\quad\quad\quad\ \neq P_{\bm{Y}}^{\mathcal{M}:\operatorname*{do}(B:=b_{1})}.\end{split}$ (4)
The proposed deduction in Eq. 4 can be tested using a paired-samples test. This hypothesis testing evaluates whether there is a significant difference between paired responses under different interventions, according to a measure of difference or similarity that is suitable for the particular context of the problem. For interval responses where a meaningful directional measure (such as scores or accuracy) can be applied, the difference is calculated using $\Delta=\bm{Y}^{\mathcal{M}:\operatorname*{do}(B:=b_{i})}-\bm{Y}^{\mathcal{M}:\operatorname*{do}(B:=b_{j})}$. In cases where the response variable is nominal and exists in a high-dimensional space without an inherent direction—such as free text responses—a similarity measure $s$ is used. This measure maps each pair of responses to a value between 0 and 1, where values closer to 1 indicate higher similarity between the responses.
When using difference for interval response, our null hypothesis $\mathcal{H}_{0}$ and alternative hypotheses $\mathcal{H}_{1}$ are
@@
When using similarity for nominal response, our null hypothesis $\mathcal{H}_{0}$ proposes that the median of the similarity exceeds a predefined threshold $\tau$, while the alternative hypothesis $\mathcal{H}_{1}$ suggests it does not.
$\mathcal{H}_{0}:m_{s}\geq\tau\text{ vs. }\mathcal{H}_{1}:m_{s}<\tau.$ (6)
@@ 2.3 Sample Generation via Indirect Intervention
Ideally, to probe the existence of IP by Eq. 4, we need to do a direct intervention on $B$ to test whether the intervened distributions are identical. However, since LLMs are complicated black-box models, directly identifying the location of $B$ and intervening on it is beyond the limit of the current interpretability research Räuker et al. (2022), not to mention the high variances of the input $\bm{X}$ and semantics $\bm{S}$, which also increase the hardness of the testing. To address these challenges, we propose a novel technique, indirect intervention, to generate approximately paired observations for the testing.
@@
The sample generation process using our indirect intervention is in Figure 2. Basically, we indirectly intervene on $B$ by generating paired observations. First, the domain $\bm{\mathcal{X}}$ of input $\bm{X}$ is divided into subspaces, each corresponding to a different $b\in\mathcal{B}$. Given a background pair $(\tilde{b}_{i},\tilde{b}_{j})$, we first generate input $\bm{x}_{i}$ from $\bm{\mathcal{X}}_{i}$, i.e., the subspace for the background $\tilde{b}_{i}$, and collect the response $\bm{y}_{i}$. Then we generate another input for the background $\tilde{b}_{j}$ by text style transfer Jin et al. (2022) to preserve the semantics. For each observed pair $\big{(}(\bm{x}_{i},\bm{y}_{i}),(\bm{x}_{j},\bm{y}_{j})\big{)}$, we set $b_{i}$ and $b_{j}$ to differ while maintaining identical hidden semantics $\bm{s}_{i}=\bm{s}_{j}$. Our sample generation process is as follows:
@@
2.4 Hypothesis Testing Method
Given a background pair $(\tilde{b}_{i},\tilde{b}_{j}),\tilde{b}_{i}\neq\tilde{b}_{j}$, a sample $\left\{\left((\bm{x}_{i}^{(k)},\bm{y}_{i}^{(k)}),(\bm{x}_{j}^{(k)},\bm{y}_{j}^{(k)})\right)\right\}_{k=1}^{n}$ is obtained using our proposed sampling method. According to the particular context of the problem, we choose difference $\Delta$ for interval response and similarity $s$ for nominal response. We test the existence of IP by testing hypothesis in Eq. 5 and Eq. 6. In our work, we set the significance level $\alpha$ to 0.05. If the derived $p$-value is less than $\alpha$, then the null hypothesis is rejected, which further implies the existence of IP. Otherwise, it means there is not enough evidence to reject $H_{0}$.
2.4.1 Permutation-based hypothesis testing for interval responses
The paired t-test, commonly used for assessing paired differences, assumes normally distributed differences Witte and Witte (2017). This assumption may not hold in cases with bounded interval responses, such as scores. To overcome this, we employ a permutation-based hypothesis testing method Good (2013), which offers greater flexibility and robustness by relaxing distributional assumptions. The test statistic and $p$-value are computed by permutation as follows:
@@ 2.4.1 Permutation-based hypothesis testing for interval responses
    Calculate the mean $\mu_{\tilde{\Delta}}^{(1)}$ of the differences after permutation, i.e., randomly reversing the sign of each difference score calculated in Step 1, and then computing the mean.
@@
    Repeat Step 3 $m$ times and collect $\{\mu_{\tilde{\Delta}}^{(i)}\}_{i=1}^{m}$.
@@
    Compute $p$-value as the proportion of permutation means that are no less than $\mu_{\Delta}$.
2.4.2 Sign-test for nominal responses
For nominal response without an inherent meaningful direction, such as free text responses. A similarity measure is used to assess the difference between responses. We test the existence of IP using hypothesis in Eq. 6. We use sign-test Manoukian (2022). Given paired responses, we first computes the similarity $\{S_{i}\}_{i=1}^{n}$. The test statistic and p-value is then computed by

### v2

@@ Abstract
Implicit Personalization (IP) is a phenomenon of language models inferring a user’s background from the implicit cues in the input prompts and tailoring the response based on this inference. While previous work has touched upon various instances of this problem, there lacks a unified framework to study this behavior. This work systematically studies IP through a rigorous mathematical formulation, a multi-perspective moral reasoning framework, and a set of case studies. Our theoretical foundation for IP relies on a structural causal model and introduces a novel method, indirect intervention, to estimate the causal effect of a mediator variable that cannot be directly intervened upon. Beyond the technical approach, we also introduce a set of moral reasoning principles based on three schools of moral philosophy to study when IP may or may not be ethically appropriate. Equipped with both mathematical and ethical insights, we present three diverse case studies illustrating the varied nature of the IP problem and offer recommendations for future research. [^1] Our code is at https://github.com/jiarui-liu/IP, and our data is at https://huggingface.co/datasets/Jerry999/ImplicitPersonalizationData.
@@ 1 Introduction
Inspired by this example, we propose the concept of Implicit Personalization (IP). Grounded in a structural causal model (SCM; Peters et al., 2017; Pearl, 2009), we define IP as a process that first infers a user’s background from the way a question is posed, and then tailors the response to fit this background, as in Figure 1. While many studies have separately explored different aspects of this problem (Flek, 2020; Raharjana et al., 2021; Eloundou et al., 2024), we still lack a community-wide standardized framework to study these phenomena. The absence of a common framework leads to divergent perspectives: some studies view it positively, suggesting that incorporating inferred user demographics can enhance NLP performance by personalized responses (Hovy, 2015; Benton et al., 2016; Sasaki et al., 2018; Salemi et al., 2024; Chen et al., 2023), whereas others criticize it negatively for introducing biases in model responses towards underrepresented groups (Bolukbasi et al., 2016; Garg et al., 2018; Arora et al., 2023; Das et al., 2023; He et al., 2024; Kantharuban et al., 2024), or for fostering flattery to satisfy users regardless of the accuracy of the information provided (Sharma et al., 2023a; Wei et al., 2023; Wang et al., 2023).
@@
RQ1: What is IP? – for which we provide a rigorous mathematical formalization (Section 2). RQ2: How to detect IP in large language models (LLMs)? – for which we provide a rigorous mathematical formulation and a set of case studies (Sections 2, 4, 5, 6 and 7). RQ3: What are the moral implications of IP? – for which we propose a framework for ethical engagement (Section 3). RQ4: How to improve future models to ensure their IP behavior aligns with ethical standards? – for which we provide a list of suggestions for developers and the community (Section 8).
By answering the above questions, our work contributes a “full-stack” systematic study on IP: In the mathematical framework, we ground IP in an SCM Peters et al. (2017); Pearl (2009), and then propose an indirect intervention method to test the causal effect in the LLM-specific, diamond-shaped causal graph in Figure 1 with un-intervenable mediators (see technical deductions in Section 2). After the technical formulation of IP, we provide a moral reasoning framework (Section 3), which connects the ethical considerations of IP to major schools of moral philosophy, including consequentialism Mill (2016); Parfit (1987), deontology Kant and Schneewind (2002); Ross (2002), and contractualism Rawls (2017); Scanlon (2000).
To illustrate the usefulness of our theoretical formulations, we present three diverse case studies that feature different instances of the IP problem (Section 4): (1) cultural adaptation, where IP is a desired model behavior (Section 5), (2) education disparity, where IP is unethical (Section 6), and (3) echo chamber, which has mixed implications for IP (Section 7). Finally, we conclude with recommendations for future research and an outlook for the community (Section 8).
@@ 2 A Mathematical Framework for IP
2.1 Causal Graph Formulation
In general, any NLP system has the functional behavior $f:\bm{x}\mapsto\bm{y}$, where the user inputs the text $\bm{x}$, and the model generates a response $\bm{y}$, as in Figure 1. Formally, IP is a sub-process within this functional behavior. In our running example, where $\bm{x}=$“What color is a football?”, the model response $\bm{y}$ should be “Brown” if IP takes place, whereas the general answer would mention both possibilities, e.g., “An American football is Brown, and a soccer ball is usually black and white.”
To evaluate the existence of IP, we model the response generation process of IP with a causal graph $\mathcal{G}$ in Figure 1. In general, a causal graph $\mathcal{G}=(\bm{V},\bm{E})$ is a good mathematical formulation for modeling cause-and-effect relationships among a set of random variables $\bm{V}$, where each edge $e_{ij}\in\bm{E}$ indicates whether the $i$-th random variable $V_{i}$ is a direct cause for the $j$-th random variable $V_{j}$ Pearl (2009); Peters et al. (2017).
In the formulated causal graph of this study in Figure 1, the language model parses $\bm{x}$, extracting its semantics $\bm{s}$—for instance, querying the color of a ball in football—and may infer the user’s background $b$ from a set of categories $\mathcal{B}$ (e.g., American English speaker). The response $\bm{y}$ is generated based on $\bm{s}$, optionally tailored to the user’s background $b$ using IP, or focusing solely on the semantics without customization.
2.2 Problem Statement
Based on our causal graph we introduce a mathematical formulation for implicit personalization. The key research question, namely “Does IP take place in LLMs?”, can be formulated as a question of causal inference, namely “Does user background ${B}$ have a causal effect on the LLM response $\bm{Y}$?”
If the cause $B$ is binary, and the effect variable $Y$ is a real number $Y\in\mathbb{R}$, this question is usually handled by estimating the average treatment effect (ATE) Pearl (1995); Pearl et al. (2000):
$\quad\mathrm{ATE}(B\rightarrow Y)$ (1)
$:=\mathbb{E}[Y|\operatorname*{do}(B=1)]-\mathbb{E}[Y|\operatorname*{do}(B=0)]~{},$ (2)
where $\operatorname*{do}(\cdot)$ denotes an intervention on the variable. The overall estimand calculates the expected change in $Y$ caused by switching $B$ from 0 to 1.
2.2.0.0.1 Quantifying the Interventional Effect: Moving from Numerical Changes to Text Changes
However, the challenge in our study is that the effect variable $\bm{Y}$ is essentially the response of an LLM in natural language, so we can no longer simplify the question to an averaged single value, but need to take into account the entire distribution $P(\bm{Y}|\operatorname*{do}(B=1))$ versus $P(\bm{Y}|\operatorname*{do}(B=0))$ under an intervention on $B$. Using the distributional change to model the causal influence was also discussed in Janzing et al. (2013) in a statistical setting. The general definition of the existence of a causal effect is as follows:
@@ 2.2.0.0.1 Quantifying the Interventional Effect: Moving from Numerical Changes to Text Changes
In the given causal graph $\mathcal{G}$, there is a causal effect from $B$ to $\mathbf{Y}$ if there exist $b_{0},b_{1}\in\mathcal{B}$, such that
$P_{\bm{Y}}^{\operatorname*{do}(B=b_{0})}\neq P_{\bm{Y}}^{\operatorname*{do}(B=b_{1})}.$ (3)
The intuition behind this is we first intervene on $B$ by setting it to different values $b_{0}$ and $b_{1}$, and then compare whether the intervened probability distributions of $\bm{Y}$ are identical. If not, then it implies that the LLM performs IP to generate different responses $\bm{Y}$ for different backgrounds $B$. Hence, our research question becomes identifying whether the pair-wise distributional change of $\bm{Y}$ is significant when perturbing $B$. We can perform the following deduction:
@@
 $&\exists b_{0},b_{1}\in\mathcal{B},\text{s.t. }P_{\bm{Y}}^{\operatorname*{do}(B=b_{0})}\neq P_{\bm{Y}}^{\operatorname*{do}(B=b_{1})}.\end{split}$ (6)
The deduction proposed in Eq. 6 can be evaluated using a paired-samples test while controlling the semantic variable $\mathbf{S}$ Witte and Witte (2017), which assesses the presence of statistically significant differences between paired responses while controlling the semantics. Depending on the characteristics of the problem, we design different statistical hypothesis tests according to the effect variable type.
First, if the effect variable is a real number, referred to as the “interval” type, the paired difference can be calculated directly using $\Delta=\bm{Y}^{\mathcal{M}:\operatorname*{do}(B:=b_{i})}-\bm{Y}^{\mathcal{M}:\operatorname*{do}(B:=b_{j})}$.
For the effect variable in high-dimensional spaces, such as free text, where subtraction is not feasible, the difference measure $\Delta$ becomes invalid. In these instances, we use a similarity metric, $s$, which assigns a value between 0 and 1 to each pair of responses, with values closer to 1 indicating higher similarity. By employing different measures for different types of responses, we transform the IP problem into a statistical hypothesis testing.
For interval responses, we want to evaluate whether there is difference between responses for different backgrounds. Our null hypothesis $\mathcal{H}_{0}$ is that there is no difference between responses for different backgrounds, i.e. the mean of the differences between the paired responses is 0. The alternative hypothesis $\mathcal{H}_{1}$ is that there is a difference in responses between different backgrounds, or that the mean of the differences is not equal to zero.
@@
For response in high-dimensional spaces where the subtraction is not applicable, we use a similarity measure $s$ to evaluate the similarity between the responses. Our null hypothesis $\mathcal{H}_{0}$ proposes that the median of the similarity is at a predefined level $M_{0}$, while the alternative hypothesis $\mathcal{H}_{1}$ suggests it’s smaller than $M_{0}$.
$\mathcal{H}_{0}:m_{s}=M_{0}\text{ vs. }\mathcal{H}_{1}:m_{s}<M_{0}.$ (8)
@@ 2.3 Sample Generation via Indirect Intervention
Ideally, to probe the existence of IP by Eq. 6, we need to do a direct intervention on $B$ to test whether the intervened distributions are identical. However, since LLMs are complicated black-box models, directly identifying the location of $B$ and intervening on it is beyond the limits of current interpretability research Räuker et al. (2022). To address these challenges, we propose a novel technique, indirect intervention, to generate approximately paired observations for testing.
@@
The sample generation process using our indirect intervention is illustrated in Figure 2. Basically, we indirectly intervene on $B$ by generating paired observations. First, the domain $\bm{\mathcal{X}}$ of input $\bm{X}$ is divided into subspaces, each corresponding to a different $b\in\mathcal{B}$. Given a background pair $(\tilde{b}_{i},\tilde{b}_{j})$, we first generate input $\bm{x}_{i}$ from $\bm{\mathcal{X}}_{i}$, i.e., the subspace for the background $\tilde{b}_{i}$, and collect the response $\bm{y}_{i}$. Then we generate another input for the background $\tilde{b}_{j}$ by text style transfer Jin et al. (2022) to preserve the semantics. For each observed pair $\big{(}(\bm{x}_{i},\bm{y}_{i}),(\bm{x}_{j},\bm{y}_{j})\big{)}$, we set $b_{i}$ and $b_{j}$ to differ while maintaining identical hidden semantics $\bm{s}_{i}=\bm{s}_{j}$. This method controls the semantics and the user’s background while randomizing hidden random variables that are not modeled in this SCM. Our sample generation process is as follows:
@@
2.4 Single Hypothesis Testing Methods
Given a pair of distinct background $(\tilde{b}_{i},\tilde{b}_{j}),\text{ where }\tilde{b}_{i}\neq\tilde{b}_{j}$, a sample $\left\{\left((\bm{x}_{i}^{(k)},\bm{y}_{i}^{(k)}),(\bm{x}_{j}^{(k)},\bm{y}_{j}^{(k)})\right)\right\}_{k=1}^{n}$ is obtained using our proposed sampling method. Depending on the type of responses, we use either the difference measure $\Delta$ for interval responses, or the similarity metric $s$ for non-interval responses, to quantify the disparity between responses. We evaluate the existence of IP by testing hypothesis in Eq. 7 or Eq. 8. In our work, we set the significance level $\alpha$ to 0.05. If the derived $p$-value is less than $\alpha$, then the null hypothesis is rejected, which further implies the existence of IP. Otherwise, it means there is not enough evidence to reject $H_{0}$.
2.4.1 Permutation-based Method for Interval Responses
The paired t-test, commonly used to assess differences between paired samples, relies on the assumption that these differences are continuous and normally distributed Witte and Witte (2017). However, this assumption does not hold in certain scenarios, such as when the data comes from rating scales. Such responses are discrete and constrained within specific bounds, violating the assumptions of the t-test. To address this limitation, we employ a permutation-based test Good (2013). This alternative does not require the normality assumption, thereby providing a more robust and flexible approach for hypothesis testing. By randomly flipping the signs of the differences between paired responses and calculating the test statistic (here the arithmetic mean) for each permutation, this method constructs an empirical distribution of the test statistic under the null hypothesis.
@@ 2.4.1 Permutation-based Method for Interval Responses
    Calculate the mean of permuted differences $\mu_{\tilde{\Delta}}^{(l)}$, i.e., randomly reversing the sign of each difference calculated in Step 1, and then compute the mean.
@@
    Repeat Step 3 $L$ times and collect $\{\mu_{\tilde{\Delta}}^{(l)}\}_{l=1}^{L}$.
@@
    Compute $p$-value as the proportion of permuted means in which the absolute value of the test statistic is greater than or equal to the absolute value of the original test statistic $\mu_{\Delta}$.
Typically, a higher number of permutations allows for a more reliable approximation of the null distribution of the test statistic. Common practice suggests using at least 1,000 permutations for general purposes, but for more rigorous studies, 10,000 or more might be necessary Good (2013).
2.4.2 Sign-Test for Non-Interval Responses
Responses from language models, such as free text, typically exist in high-dimensional spaces where subtraction $\Delta$ is invalid to measure the disparity between these responses. To study how the responses vary across different backgrounds, we introduce a similarity measure, which quantifies the difference between pairs of responses on a scale from 0 to 1, with values closer to 1 indicating higher similarity.
To evaluate the existence of IP, we test whether the median of the similarity scores is smaller than a predefined threshold $M_{0}$ or not, as in Eq. 8. Since the similarity measure does not conform to the assumptions required for a t-test, namely normal distribution, a non-parametric alternative, sign-test Conover (1999), is used. The sign test evaluates the hypothesis by counting how many data points are below the specified median, assuming that under the null hypothesis, each data point is equally likely to be above or below the median, which can be formulated using a binomial distribution. The p-value is then computed as the probability of obtaining test statistic at least as extreme as the observed one.